OpenStudy (anonymous):

Solve x^2 – 5x + 10 = 0 by completing the square.

OpenStudy (anonymous):

whr r u stuck?

OpenStudy (anonymous):

I'm stuck on the beginning

OpenStudy (anonymous):

\[\left( x-a \right)^{2} = x^{2} + a^{2} - 2ax\]

OpenStudy (anonymous):

x2+a2−2ax arrange it like this

OpenStudy (anonymous):

x^2 + 10 = 5x?

OpenStudy (anonymous):

x + root10 will be sqaured

OpenStudy (anonymous):

ok so can it be (x - 5/2)^2 + 15/4 = 0?

OpenStudy (anonymous):

no \[\left( x+\sqrt{10} \right)^{2}\]